-- SebastianRabien - 12 Aug 2016

Julian and SR on the mountain. Actions performed:
  • The new remote power is installed and tested. New engineering gui installed by Jose. Controiller not yet usable with it!!!
  • The power analyzer can now measure UPS current,
  • The instrument alarm panel runs permanently on UPS,
  • JZ measured the propagation shutter speed (results later),
  • The LLP SX rack is prepared for the TBAD delay installation
  • we removed a big heap of junk from the ARGOS cabinets.
  • BNC cables for the BCU-Piezo control NOT routed yet
  • measured the first pulse properties for all lasers. look all more or less like that:
  • means the FPK is ok for a the lasers. (top trace are the laser pulses, middle: trigger, low: burst trigger)
  • We read out all innolas laser parameters from the units for reference and cross check.
Topic revision: r1 - 12 Aug 2016, SebastianRabien
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback